OpenStudy (anonymous):

Which shows 54^2 − 46^2 being evaluated using the difference of squares method? 54^2 − 46^2 = (2916 + 2116)(2916 − 2116) = 4,025,600 54^2 − 46^2 = (54 + 46)(54 − 46) = (100)(8) = 800 54^2 − 46^2 = 2916 − 2116 = 800 54^2 − 46^2 = (54 − 46)^2 = 8^2 = 64

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

using letters difference of 2 squares is a^2 - b^2 = (a + b)(a - b) same 2 letters in each bracket but + in one bracket and - in the other

OpenStudy (welshfella):

so if you put a = 54 and b = 46 in the above you can see which choice is correct

OpenStudy (welshfella):

54^2 - 46^2 = ( 54 + )( 54 - ) = ? can you fill in the blanks?

OpenStudy (anonymous):

46?

OpenStudy (welshfella):

right now you should know the correct choice

OpenStudy (anonymous):

Yep :]

OpenStudy (anonymous):

B correct?

OpenStudy (welshfella):

yes